OpenStudy (anonymous):

Solve for X: (3^2x-1) - 19(3^x) + 20 = 0

Parth (parthkohli):

\[\dfrac{1}{3}\cdot3^{2x} - 19 \cdot 3^x + 20 = 0\]If you let \(t = 3^x\), then\[\dfrac{1}{3}t^2 - 19 t + 20 = 0\]

OpenStudy (anonymous):

why the first term be 1/3?

Parth (parthkohli):

\[3^{2x-1} = \dfrac{3^{2x}}{3^1} = \frac{1}{3}\cdot3^{2x}\]

OpenStudy (anonymous):

Ohh thank you soo much!

Parth (parthkohli):

No problem!

OpenStudy (anonymous):

@ParthKohli Would you have to use the quadratic formula to find t then set it equal to 3x?

Parth (parthkohli):

Exactly. Ignore negative solutions, if any.

OpenStudy (anonymous):

@ParthKohli So far I have \[1/2 ( 57+\sqrt{3009}) =3x\]

Parth (parthkohli):

Yeah, that is one solution. But you shouldn't ignore the other one.

OpenStudy (anonymous):

right then just divide both by 3?